OpenStudy (anonymous):

how to prove, sin^-1 x = tan^-1 x/√(1-x^2)

OpenStudy (zarkon):

let \[\theta=\sin^{-1}(x)\] then \[\sin(\theta)=x\]  thus \[\tan(\theta)=\frac{opp}{adj}=\frac{x}{\sqrt{1-x^2}}\]